OpenStudy (anonymous):

3 = log8 + 3logx need help please and thankyou

OpenStudy (anonymous):

start by using \(n\log(x)=\log(x^n)\) to rewrite it as \[3=\log(8)+\log(x^3)\]

OpenStudy (anonymous):

then use \(\log(A)+\log(B)=\log(AB)\) to rewrite it as \[3=\log(8x^3)\]

mathslover (mathslover):

\(\log a + \log b = \log (ab) \) - Next step

OpenStudy (callisto):

log 8 = log (2^3) = 3 log 2 3 = log8 + 3logx 3 = 3 log 2 + 3 log x 1 = log 2 + log x

mathslover (mathslover):

here, a = 8 and b = x^3 ...

OpenStudy (anonymous):

then rewrite in equivalent exponential form assuming that the log is log base ten, exponential form is \[10^3=8x^3\]

OpenStudy (anonymous):

then your puttin 3=x+3

mathslover (mathslover):

Next step : \(3 = \log (2x)^3 \) \(3 = 3 \log (2x) \) \(1 = log (2x)\)

OpenStudy (anonymous):

solve for \(x\) by dividing by \(8\) and then taking the cubed root of both sides

mathslover (mathslover):

That is : \(1 = \log (2x) \) If base is 10, then 2x = 10 as : \(log_ a a = 1\)

OpenStudy (anonymous):

x=5

mathslover (mathslover):

solve it for x further. \(\bf{2x = 10}\)

OpenStudy (anonymous):

\[10^3=8x^3\] \[\frac{10^3}{8}=x^3\] \[\frac{10}{2}=x\]\[5=x\]

mathslover (mathslover):

Simply use : \(\log _ a a = 1\) , if base is 10, then 2x has to be 10 also. That's it...